OpenStudy (anonymous):

Write an equation of a line that is perpendicular to y = 6x -9 and has a y-intercept of 3.

OpenStudy (anonymous):

You can find the gradient of perpendicular line using the formula M1m2=-1

OpenStudy (anonymous):

no i dont get this stuff

OpenStudy (anonymous):

Okay the general equation of a line is y=mx+c where m is your gradient/slope and c is your y-intercept

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Okay so for you to find the perpendicular line, u need to find y=mx+c Do you know your C?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

Good, so now u just have to find your m. There is a formula that says that the gradient of the first equation X gradient of perpendicular equation = -1 Hence M1M2=-1 From there find M2

OpenStudy (anonymous):

i am not sure i under stand your last reply?

OpenStudy (anonymous):

what do i do?

OpenStudy (anonymous):

Okay so basically need to find the gradient of the perpendicular line.

OpenStudy (anonymous):

y=mx+c to y=m(-1)+3??

OpenStudy (anonymous):

No, to find the gradient of your perpendicular line : Gradient of original line X Gradient of perpendicular line =-1 M1*m2 =-1 6*m2 =-1 m2 = -1/6

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

:) Thank you!